# Global identification of dynamic panel models with interactive effects

### v1

@@ 5 Individual fixed-effects
Cases 1’ and 2’ correspond to the case in which $\widetilde{J}_{(1,2,3),(4,5,6)}$ is the 0-polynomial, so that all the coefficients are trivially equal to 0. Cases 3’ and 4’ correspond to the case in which we can found a root $\widetilde{\alpha}^{\star}$ that solves $\widetilde{J}_{(1,2,3),(4,5,6)}(\widetilde{\alpha}^{\star},\theta)=0$. However, once $\widetilde{\alpha}$ is pinned-down in this way, we have no other free parameter, so the same root must satisfy all the 0-determinant conditions (Claim 1) associated with the remaining diagonal exclusion minors of $\mathbf{O}$. Case 5’ corresponds to the case where global identification is achieved.
@@
This case implies that $\det(\mathbf{\Psi})=0$, which is analogous to the case where $\Psi=0$ in subsection 4.3. However, this case is excluded because we assume that $\mathbf{\Psi}$ is positive definite.
@@
From Assumptions 4’ and 5’, we know that $d_{1},d_{2},d_{3},\alpha$ are fixed population parameters, while $f_{1},f_{2},f_{3},f_{\gamma}$ are drawn from a continuous probability distribution. Therefore, we can apply Lemma 2 to conclude that Case 2’ occurs only with probability zero.
@@
We demonstrate that the case $\widetilde{\alpha}=1$ can never satisfy 5 with positive probability. The proof for $\widetilde{\alpha}=-\frac{1}{f_{4}}$ is analogous and thus omitted.
If we set $\widetilde{\alpha}=1$ in order to satisfy Equation 20, then we have no other free parameters. Thus, this same $\widetilde{\alpha}$ must satisfy all the 0-determinant conditions (Claim 1). However, this cannot occur with positive probability.
To see this, consider the determinant of the diagonal exclusion minor associated with rows $R=(1,5,6)$ and columns $C=(2,3,4)$. By Proposition 1 and Corollary 3, we know it can be written as:
$\det\left(\mathbf{M}^{\mathbf{O},3}_{(1,5,6),(2,3,4)}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta\right)$
where $\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta\right)$ is a random polynomial in $\widetilde{\alpha}$ of order 4. Case 4’ is plausible if both $\widetilde{J}_{(1,2,3),(4,5,6)}\left(\widetilde{\alpha},\theta\right)$ and $\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta\right)$ share the root $\widetilde{\alpha}=1$. Note that this is equivalent to stating that $\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta\right)|_{\widetilde{\alpha}=1}=0$. Evaluating $\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta\right)$ at $\widetilde{\alpha}=1$ yields:
$\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta\right)|_{\widetilde{\alpha}=1}$ $=\left(d_{4}f_{2}-d_{4}f_{3}-\alpha d_{3}f_{2}+\alpha d_{3}f_{4}+\alpha^{2}d_{2}f_{3}-\alpha^{2}d_{2}f_{4}\right)$
×$\quad\times\left(f_{\gamma}^{2}\Psi_{1,2}^{2}-f_{\gamma}\Psi_{1,2}^{2}-d_{1}\Psi_{2,2}-f_{\gamma}^{2}\Psi_{1,1}\Psi_{2,2}\right.$
@@
Thus, Case 4’ is plausible if either term is equal to zero. However, this can never occur with positive probability. Let’s begin with the first term. Using Assumption 6’, the case in which this term equals zero can be rewritten as:
$f_{2}=\frac{(d_{4}-\alpha^{2}d_{2})f_{3}+(\alpha^{2}d_{2}-\alpha d_{3})f_{4}}{d_{4}-\alpha d_{3}}$
This equation defines a plane on which the three-dimensional random variable $(f_{2},f_{3},f_{4})$ must lie. Since this is a probability zero event (Lemma 2), it cannot occur with positive probability. Now, consider the second term. It can be expressed as a second-order polynomial in $f_{\gamma}$:
@@
Setting the second term equal to zero is equivalent to finding the roots of this polynomial. First, note that $\mathcal{P}$ can never be the zero polynomial since $b>0$, which holds because $\mathbf{\Psi}$ is positive definite by assumption. Then, by the fundamental theorem of algebra, there are at most two values of $f_{\gamma}$ that satisfy $\mathcal{P}(f_{\gamma})=0$. Since $f_{\gamma}$ is assumed to be drawn from a continuous distribution, the event $\mathcal{P}(f_{\gamma})=0$ has zero probability, because these two roots have zero Lebesgue measure in the support of $f_{\gamma}$.
Thus, we can conclude that almost surely, Case 4’ is not plausible. A similar argument leads to the conclusion that Case 3’ can only occur with probability zero.
We have shown that Case 5’ is the only alternative that makes $\det\left(\mathbf{M}^{\mathbf{O},3}_{(1,2,3),(4,5,6)}\right)=0$ with positive probability. Thus, $\alpha$ is almost surely globally identified in the presence of individual fixed effects.
@@
Our entire proof carries over without any modification. $\Box$

### v2

@@ 5.1 Analysis in levels
Cases 1’ and 2’ correspond to the case in which $\widetilde{J}_{(1,2,3),(4,5,6)}$ is the zero polynomial, so that all the coefficients are trivially equal to 0. Cases 3’ and 4’ correspond to the case in which we can find a root $\widetilde{\alpha}^{\star}$ that solves $\widetilde{J}_{(1,2,3),(4,5,6)}\left(\widetilde{\alpha}^{\star},\theta^{0}\right)=0$. However, once $\widetilde{\alpha}$ is pinned-down in this way, we have no other free parameter, so the same root must satisfy all the zero-determinant conditions (Claim 1) associated with the remaining diagonal exclusion minors of $\mathbf{O}$. Case 5’ corresponds to the case where global identification is achieved.
@@
This implies that $\det(\mathbf{\Psi})=0$, which is analogous to the case where $\Psi=0$ in subsection 4.3. However, this possibility is excluded because we assume that $\mathbf{\Psi}$ is PD.
@@
From Assumptions B6 and B7, we know that $d_{1},d_{2},d_{3},\alpha$ are fixed population parameters, while $f_{1},f_{2},f_{3},f_{\gamma}$ are drawn from a continuous probability distribution. Therefore, we can apply Lemma 2 to conclude that Case 2’ occurs only with probability zero. Note that given that $d_{3}>0$ then the expression cannot be trivially equal to 0.
@@
We demonstrate that the case $\widetilde{\alpha}=1$ can never satisfy (5) with positive probability. The proof for $\widetilde{\alpha}=-\frac{1}{f_{4}}$ is analogous and thus omitted. If we set $\widetilde{\alpha}=1$ in order to satisfy (22), then we have no other free parameters. Thus, this same $\widetilde{\alpha}$ must satisfy all the zero-determinant conditions (Claim 1). However, this cannot occur with positive probability. To see this, consider the determinant of the diagonal exclusion minor associated with rows $R=(1,5,6)$ and columns $C=(2,3,4)$. By Proposition 1 and Corollary 3, we know it can be written as:
$\det\left(\mathbf{M}^{\mathbf{O},3}_{(1,5,6),(2,3,4)}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta^{0}\right)$
where $\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta^{0}\right)$ is a random polynomial in $\widetilde{\alpha}$ of degree 4. Case 4’ is plausible if both $\widetilde{J}_{(1,2,3),(4,5,6)}\left(\widetilde{\alpha},\theta^{0}\right)$ and $\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta^{0}\right)$ share the root $\widetilde{\alpha}=1$. Note that this is equivalent to stating that $\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta^{0}\right)|_{\widetilde{\alpha}=1}=0$. Evaluating $\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta^{0}\right)$ at $\widetilde{\alpha}=1$ yields:
$\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta^{0}\right)|_{\widetilde{\alpha}=1}$ $=\left(d_{4}f_{2}-d_{4}f_{3}-\alpha d_{3}f_{2}+\alpha d_{3}f_{4}+\alpha^{2}d_{2}f_{3}-\alpha^{2}d_{2}f_{4}\right)$
$\quad\times\left(f_{\gamma}^{2}\Psi_{1,2}^{2}-f_{\gamma}\Psi_{1,2}^{2}-d_{1}\Psi_{2,2}-f_{\gamma}^{2}\Psi_{1,1}\Psi_{2,2}\right$.
@@
Thus, Case 4’ is plausible if either term is equal to zero. However, this can never occur with positive probability. Let’s begin with the first term. The case in which this term equals zero can be rewritten as:
$(d_{4}-\alpha d_{3})f_{2}-(d_{4}-\alpha^{2}d_{2})f_{3}-(\alpha^{2}d_{2}-\alpha d_{3})f_{4}=0$
This equation defines a plane on which the three-dimensional random variable $(f_{2},f_{3},f_{4})$ must lie. By Lemma 2 and Assumption B8, this is a zero probability event. Now, consider the second term. It can be expressed as a second-degree polynomial in $f_{\gamma}$:
@@
Setting $\mathcal{P}(f_{\gamma})$ equal to zero is equivalent to finding the roots of this polynomial. First, note that $\mathcal{P}$ can never be the zero polynomial since $b>0$, which holds because $\mathbf{\Psi}$ is PD by assumption. Then, by the fundamental theorem of algebra, there are at most two real values of $f_{\gamma}$ that satisfy $\mathcal{P}(f_{\gamma})=0$. Since $f_{\gamma}$ is assumed to be drawn from a continuous distribution, the event $\mathcal{P}(f_{\gamma})=0$ has zero probability, because these two roots have zero Lebesgue measure in the support of $f_{\gamma}$.
Thus, we can conclude that almost surely, Case 4’ is not plausible. A similar argument leads to the conclusion that Case 3’ can only occur with probability zero. We have shown that Case 5’ is the only alternative that makes $\det\left(\mathbf{M}^{\mathbf{O},3}_{(1,2,3),(4,5,6)}\right)=0$ with positive probability. This implies that $\widetilde{\alpha}=\alpha$ is a necessary condition for (5) to hold in the presence of individual fixed effects. Thus, $\alpha$ is almost surely globally identified.
@@
Our entire proof carries over without any modification. $\square$